The last time Jim went fishing he only caught 3 fish. The total weight was 64 ounces. The first one was twice as big as the second fish. And the third one was halve as large as the first. How many ounces did each fish weigh?
Answer by 303795(562)

Call the size of the second fish x.
The first fish is twice as big so it is 2x.
The third fish is half the size of the first fish so it is 1/2 of 2x or just x.
The weight of the three fish added together is 2x+x+x=4x which is 64 ounces.
so 4x=64
x=64/4=16
So the first fish had a weight of 32 ounces and the other two fish each weighed 16 ounces.